# TaskCLIP: Extend Large Vision-Language Model for Task Oriented Object Detection

### v1

@@ 4.5 Visualization and Discussion
Fig. 5 presents our model’s predictions alongside the ground truth on various image samples. Specifically, Fig. 5 (a) showcases examples where our model performs well. The predicted bounding boxes exhibit great overlap with the ground truth across diverse tasks, demonstrating the robust generalizability of our model.
However, Fig. 5 (b) highlights two examples where our model’s performance is suboptimal. These examples illustrate two common challenges faced by our model. Firstly, in the upper panel of Fig. 5 (b), there are two chairs depicted, but only the distant chair can be utilized to open a bottle of beer due to a person occupying the nearby chair. While the ground truth accurately captures this subtle distinction, our model erroneously identifies all objects resembling chairs based solely on visual features. Additionally, in the lower panel of Fig. 5 (b), our model erroneously selects both a cup and a tie to extinguish fire. This misprediction arises from the tie’s visual features closely resembling those provided by LLM for solving this task, such as a sturdy handle and a bucket shape. Consequently, our model mistakenly identifies the tie due to its misleading appearance.
@@ 5 Conclusion
In this study, we introduce TaskCLIP, a novel framework for task-oriented object detection. TaskCLIP leverages pre-trained knowledge and vision-language associations from the frozen CLIP model in an efficient manner, distinguishing itself from prior research efforts. To address the challenge of misalignment between object image embeddings and their corresponding visual attributes, we propose a transformer-based aligner to recalibrate the vision and text embedding space. Additionally, we propose a select-by-grouping mechanism to mitigate the issue of high false negative mispredictions stemming from imbalanced training data. This mechanism efficiently utilizes the classification output of the object detection network. We validate the efficacy of TaskCLIP through empirical experiments, achieving SOTA performance on the COCO-Tasks dataset. Comparative analysis with prior DETR-based approaches demonstrates TaskCLIP’s superiority in terms of both accuracy and training efficiency.

### v2

@@ 4.5 Visualization and Discussion
Fig. 4 presents the predictions of our model alongside the ground truth in various image samples. Specifically, Fig. 4 (a) showcases examples where our model performs well. The pridictions exhibit great overlap with the ground truth across various tasks, demonstrating the robust generalizability of our model. However, Fig. 4 (b) highlights two examples where our model’s performance is sub-optimal. These examples illustrate two common challenges faced by our model. Firstly, in the upper panel of Fig. 4 (b), there are two chairs depicted, but only the distant chair can be utilized to open a bottle of beer due to a person occupying the nearby chair. While the ground truth accurately captures this subtle distinction, our model erroneously identifies all objects resembling chairs based solely on visual features. Furthermore, in the lower panel of Fig. 4 (b), our model erroneously selects both a cup and a tie to extinguish fire. This misprediction arises from the tie’s visual features closely resembling those provided by LLM for solving this task, such as a sturdy handle and a bucket shape. Consequently, our model mistakenly identifies the tie due to its misleading appearance.
@@ 5 Conclusion
In this study, we introduce TaskCLIP, a novel framework for task-oriented object detection. TaskCLIP leverages pre-trained knowledge and vision-language associations from the frozen CLIP model in an efficient manner, distinguishing itself from prior research efforts. Comparative analysis with prior DETR-based approaches demonstrates TaskCLIP’s superiority in terms of both task generability, accuracy, and training efficiency.